and y components of the electric field pro- 17-53 at (a) point A; Find the (b) point B. 91 +2.00 μC 42=-1.00 μC y, m d2 Fig. 17-53 2- 91
Answer #1

According to the given problem,

For point A Ex comes from q1 [positve]and is directed to the right

So Ex = k*q1/r2 = 9.0*109*2*10-6/22 = 4500N/C

Ey comes from q2 [negative]and is directed upward

So Ey = k*q2/r2 = 9.0*109*1*10-6/22 = 2250N/C

For point B:

Ex comes from both q1 & q2

from q1 the x field is to the right

Ex1 = k*q1/r2 = 9*109*2*10-6/42 = 1125 N/C

q2 contributes both an x & y component

r for this charge is (22 + 22) = 2.828m and the angle it makes with the x axis is 135o (since the field is toward q2

so Ex2 = k*q2/r2*cos(θ) = 9.0*109*1*10-6/2.8282*cos(135o) = -795.5 N/C

and Ey2 = k*q2/r2*sin(θ) = 9.0*109*1*10-6/2.8282*sin(135o) = 795.5 N/C

So Ex = Ex1 + Ex2 = 329.5N/C

and Ey = 795.5N/C